# Larger cities, more commuters, more crime? The role of inter-city commuting in the scaling of urban crime

### v1

@@ Abstract.
Cities attract a daily influx of non-resident commuters, reflecting their role in wider urban networks—not as isolated places. However, it remains unclear how this interconnectivity shapes the way crime scales with population, given that larger cities tend to receive more commuters and experience more crime. Here, we investigate how inter-city commuting relates to the population–crime relationship. We find that larger cities receive proportionately more commuters, which in turn is associated with higher crime levels. Specifically, each 1% increase in inbound commuters corresponds to a 0.32% rise in theft and 0.20% rise in burglary, holding population constant. We show that models incorporating both population and commuter inflows better explain crime variation than population-only models. These findings underscore the importance of considering how cities are connected—not just their population size—in disentangling the population–crime relationship.
@@ Introduction
Rather than isolated entities, cities are part of networks of exchange—of goods, people, and ideas—and attract outsiders through economic and social opportunities. Early works have recognized that “singular cities” concentrate inflows from their surroundings, drawing in individuals who may face increased risks of offending or victimization (?, ?, ?, ?). This inter-city dependence means that city-level crime analyses ignoring meso- or micro-level relationships risk producing bias (?, ?, ?). Likewise, daily inflows in cities create a discrepancy between residential and ambient populations, and neglecting them can distort crime rates (?). To address this distortion, previous works have suggested estimating ambient population and accounting for the number of targets (?, ?, ?). Yet, despite these methodological advances, prior studies have failed to understand whether and how inter-city connectivity itself relates to the relationship between population size and crime.
@@
Here, we investigate how inter-city commuting relates to the population–crime relationship. We use data on 8 million workers in England and Wales to construct a commuting network and assess how crime scales with population and inbound commuters. We find that inbound commuting grows proportionally with city size, and that models including both population and commuter inflows provide a more parsimonious and better-fitting explanation of crime levels than those based on population alone. Crucially, theft and burglary scale superlinearly with the combined effect of population and commuting—but once commuter inflows are accounted for, the relationship between crime and population becomes approximately linear. We find that each 1% increase in inbound commuters corresponds to a 0.32% rise in theft and 0.20% in burglary, holding population constant. Our work demonstrates that inter-city connectivity is a crucial factor in crime scaling, revealing that disentangling the population–crime relationship requires considering how cities are connected to one another.
@@ Results
Disentangling the role of inbound commuters in the scaling of crime
To set a baseline for understanding how commuter flows relate to crime, we first establish how crime scales with population size, replicating the population–crime relationship reported in previous studies. We estimate the scaling exponent $\beta$ from $C\sim N^{\beta}$, where $C$ is the number of crimes and $N$ is the population size, focusing on theft and burglary (see Methods). Our estimates align with prior findings for the UK (?). We find that both crimes scale superlinearly with population, with $\beta=1.22$ (95% CI: 1.14–1.29) for theft and $\beta=1.25$ (95% CI: 1.18–1.32) for burglary (Fig. 2ab). This superlinearity implies that a 1% increase in population is associated with more than a 1% increase in crime—a 1% population growth corresponds to a 1.22% increase in theft and 1.25% in burglary.
@@
Larger cities have more crime—and more commuters. Next, we disentangle how crime scales with both population and inbound commuters (Fig. 2cd). To do so, we extend our model to include both resident population and inbound commuters. We model crime as a function of population and commuter inflows, $C\sim N^{\gamma}I^{\alpha}$, where $N$ is the population size and $I$ is the number of inbound commuters. This formulation follows the Cobb–Douglas production function, commonly used to capture the joint contribution of multiple factors (?, ?). We estimate the exponents $\gamma$ and $\alpha$ from the data and compare this model against the population-only version using the Bayesian Information Criterion (BIC). We find that the population-and-commuter model provides a more parsimonious and accurate fit. For theft, we find $\gamma=0.92$ (95% CI: [0.83, 1.01]) and $\alpha=0.32$ (95% CI: [0.25, 0.39]); for burglary, $\gamma=1.06$ (95% CI: [0.97, 1.15]) and $\alpha=0.20$ (95% CI: [0.14, 0.27]) (see Fig. 2ef). These estimates, along with the better fit of the model relative to the population-only version, suggest that population and inter-city commuting jointly contribute to the observed scaling of urban crime.
@@ Discussion
We find that models including both population size and inter-city commuting explain crime variations more accurately and parsimoniously than population-only models. We show that, though theft and burglary scale superlinearly with population, the relationship becomes linear once we adjust for inbound commuters. When we isolate the effect of commuting, we find that higher inbound flows are associated with increased crime.
These findings contribute to prior research that underscores the need to consider both local conditions and inter-city dynamics to analyze crime in cities, including early works proposing that cities’ regional dominance should be considered when evaluating crime rates (?, ?) and more recent evidence that neglecting meso-level processes can bias city-level analyses (?). Our results reinforce the view that population-only models overlook the broader urban context. By combining scaling analysis with commuting data, we explicitly incorporate inter-city connectivity and quantify how commuter inflows shape the expected scaling of crime with population size.
@@ Methods
To explore the relationship between crime and commuting, we use police-recorded crime data from 2011/2012 to match the census period. Recording practices vary across police forces and between countries in the UK. In England and Wales, crime data is attributed to Community Safety Partnerships (CSPs), local collaborations that aim to reduce crime and improve community safety. For our analysis, we use the “Police Recorded Crime Community Safety Partnership Open Data Table, year ending March 2012” data set, which provides detailed records of all recorded offences per CSP from April 2011 to March 2012  (?). Because this data set is more detailed and consistent than those available for other parts of the UK, we focus exclusively on England and Wales, excluding Scotland and Northern Ireland from our crime analysis. Additionally, since some CSPs encompass multiple local authorities—for example, North Yorkshire CSP includes Craven, Hambleton, Harrogate, Richmondshire, Ryedale, Scarborough, and Selby—we pre-processed the data to aggregate nodes from the local authority commuting data network to their corresponding CSPs used in the 2011 crime data set. To associate each LAD with its respective CSP, we use the “LAD to Community Safety Partnership to PFA (December 2019)” data set (?), which provides the most consistent structural match with the 2011 Census commuting data. The “UK Local Authorities (past and current)” data set (?) was used to obtain LAD codes—essential for linking LADs to CSPs—and to retrieve geographic coordinates for constructing the England and Wales network. Finally, we used the “2011 Census: Usual resident population, local authorities in England and Wales” ’ to obtain the estimated total population for each LAD (?), which we preprocessed by aggregating multiple LADs into their corresponding CSPs and summing their populations where CSPs spanned more than one LAD.
@@
For burglary, Model 2 offers the best fit based on BIC and includes only statistically significant predictors (Table 1). For theft, Model 3 slightly outperforms Model 2 in BIC but only marginally; given the minimal gain in explanatory power and added complexity, Model 2 remains the most robust and interpretable choice (Table 2).
@@
BIC $-328.3$ $-357.6$ $-354.9$
F-statistic $1317.0$*** $750.8$*** $504.7$***
Durbin-Watson $2.15$ $2.17$ $2.16$
Cond. No. $121$ $158$ $6470$

### v2

@@ Abstract.
Cities attract a daily influx of non-resident commuters, reflecting their roles within wider urban networks—not as isolated places. However, it remains unclear how this interconnectivity shapes the way crime scales with population, given that larger cities tend to receive more commuters and experience more crime. In this work, we investigate how inter-city commuting relates to the population–crime relationship. We find that larger cities receive proportionately more commuters, which in turn is associated with higher levels of burglary, drug possession, robbery, shoplifting, and theft. For example, each 1% increase in inbound commuters corresponds to a 0.32% rise in theft and 0.20% rise in burglary, holding population size constant. We demonstrate that models incorporating both population size and commuter inflows explain variation in these offenses better than population-only models. Our findings underscore the importance of considering how cities are connected—not just their population size—in disentangling the population–crime relationship.
@@ Introduction
Rather than isolated entities, cities are part of networks of exchange—of goods, people, and ideas—and attract outsiders through economic and social opportunities. Early works have recognized that “singular cities” concentrate inflows from their surroundings, drawing in individuals who may face increased risks of offending or victimization (?, ?, ?, ?). This inter-city dependence means that city-level crime analyses ignoring meso- or micro-level relationships risk producing bias (?, ?, ?). Likewise, daily inflows in cities create a discrepancy between residential and ambient populations, and neglecting them can distort crime rates (?), a mismatch also evident at finer scales where crime scales with floating populations (?). To address this distortion, previous works have suggested estimating ambient population and accounting for the number of targets (?, ?, ?). Yet, despite these methodological advances, prior studies have failed to understand whether and how inter-city connectivity itself relates to the relationship between population size and crime.
@@
Here, we investigate how inter-city commuting relates to the population–crime relationship. We use data on 8 million workers in England and Wales to construct a commuting network and assess how crime scales with population and inbound commuters. We find that inbound commuting grows proportionally with city size, and that models incorporating both population and commuter inflows provide a more parsimonious and better-fitting explanation of crime levels than models based on population alone for burglary, drug possession, robbery, shoplifting, and theft. In contrast, drug trafficking and homicide show no evidence of association with inter-city connectivity. For the offenses associated to commuting, we find that crime levels scale superlinearly with the combined effect of population and commuting—but once commuter inflows are accounted for, the relationship between crime and population becomes approximately linear. Our work demonstrates that inter-city connectivity is a crucial factor in crime scaling, revealing that disentangling the population–crime relationship requires considering how cities are connected to one another.
@@ Results
Larger cities, more commuters, more crime
To set a baseline for understanding how commuter flows relate to crime, we first establish how crime scales with population size, replicating the population–crime relationship reported in previous studies. We estimate the scaling exponent $\beta$ from $C\sim N^{\beta}$, where $C$ is the number of crimes and $N$ is the population size, focusing on theft and burglary first, before turning to other crime types (see Methods). Our estimates align with prior findings for the UK (?). We find that both crimes scale superlinearly with population, with $\beta=1.22$ (95% CI: 1.14–1.29) for theft and $\beta=1.25$ (95% CI: 1.18–1.32) for burglary (Fig. 2ab). This superlinearity implies that a 1% increase in population is associated with more than a 1% increase in crime—a 1% population growth corresponds to a 1.22% increase in theft and 1.25% in burglary.
@@
Larger cities have more crime—and more commuters. Next, we disentangle how crime scales with both population and inbound commuters (Fig. 2cd). To do so, we extend our model to include both resident population and inbound commuters. We model crime as a function of population and commuter inflows, $C\sim N^{\gamma}I^{\alpha}$, where $N$ is the population size and $I$ is the number of inbound commuters. This formulation follows the Cobb–Douglas production function, commonly used to capture the joint contribution of multiple factors (?, ?), and builds on prior applications of production-function models to urban output (?, ?). We estimate the exponents $\gamma$ and $\alpha$ from the data and compare this model against the population-only version using the Bayesian Information Criterion (BIC). We find that the population-and-commuter model provides a more parsimonious and accurate fit. For theft, we find $\gamma=0.92$ (95% CI: [0.83, 1.01]) and $\alpha=0.32$ (95% CI: [0.25, 0.39]); for burglary, $\gamma=1.06$ (95% CI: [0.97, 1.15]) and $\alpha=0.20$ (95% CI: [0.14, 0.27]) (see Fig. 2ef). These estimates, along with the better fit of the model relative to the population-only version, suggest that population and inter-city commuting jointly contribute to the observed scaling of urban crime.
@@
Next, we extend our analysis to other crime types—drug possession, drug trafficking, homicide, robbery, and shoplifting—to examine how the influence of commuting varies across offenses. Our results reveal that the population-and-commuter model outperforms the population-only model for drug possession, robbery, and shoplifting; in contrast, the population-only model provides a better fit for drug trafficking and homicide (see Fig. 3). We find $\gamma=1.05$ (95% CI [0.91, 1.20]) and $\alpha=0.25$ (95% CI [0.15, 0.36]) for drug possession; $\gamma=1.25$ (95% CI [0.99, 1.50]) and $\alpha=0.89$ (95% CI [0.70, 1.08]) for robbery; and $\gamma=1.08$ (95% CI [0.94, 1.22]) and $\alpha=0.16$ (95% CI [0.05, 0.26]) for shoplifting (see Methods). Similar to theft and burglary, these offenses scale superlinearly with population but become approximately linear once commuting is included. Notably, drug trafficking and possession diverge: though both scale superlinearly with population ($\beta\approx 1.3$), only possession is associated with inbound commuting, highlighting their distinct underlying mechanisms, otherwise hidden by population-only scaling.
Figure 3: The role of inter-city commuting across crime types. We estimate population and inbound commuter coefficients for each offense using the population-and-commuter model. This model explains variation in burglary, drug possession, robbery, shoplifting, and theft; in contrast, drug trafficking and homicide show no evidence of association with inbound commuting. The asterisks indicate where the population-and-commuter model offers a better fit than the population-only specification, whereas the bars denote 95% confidence intervals.
@@ Discussion
We find that models including both population size and inter-city commuting explain crime variations more accurately and parsimoniously than population-only models in the case of burglary, drug possession, robbery, shoplifting, and theft. We show that, though these offenses scale superlinearly with population, the relationship becomes linear once we adjust for inbound commuters. When we isolate the effect of commuting, we find that higher inbound flows are associated with increased crime.
In contrast, we find no evidence that inbound commuting is associated with homicide or drug trafficking. In particular, it is noteworthy that drug trafficking and drug possession yield distinct results. Though both scale superlinearly with population, only possession is associated with inbound commuting, bringing out their distinct underlying mechanisms. This distinction would remain unobserved in traditional population-only scaling analyses, showing that superlinearity can arise with or without inter-city connectivity and demonstrating the need to account for connectivity when examining the population–crime relationship.
Our findings support early ecological perspectives that view cities as interdependent rather than isolated units, where dominant centers attract nonresidents and expand the pool of potential offenders and victims beyond the resident population (?, ?). These studies argued that variation in city crime rates reflects ecological position and dominance within the metropolitan system rather than solely local social or economic conditions. Our results provide quantitative evidence for this view, suggesting that analyses of urban crime must account for both local conditions and inter-city dynamics. They also corroborate recent evidence that neglecting meso-level processes can bias city-level analyses (?). Together, these findings reinforce that population-only models miss the broader urban context. By combining scaling analysis with commuting data, we explicitly incorporate inter-city connectivity and quantify how commuter inflows shape the scaling of crime with population size.
These results also have broader implications for the theories of urban scaling. Most mechanistic models describe superlinear growth as the outcome of processes occurring within cities, such as intensified social interactions and increasing functional diversity (?), with a few notable exceptions that incorporate inter-city processes (?). Our results show that inter-city connectivity contributes to scaling patterns, corroborating recent evidence of spatial and network effects linking urban systems (?, ?). The association between crime and commuter inflows indicates that processes shaping urban outcomes operate across, not only within, city boundaries, suggesting that mechanistic models of crime should account for inter-city connectivity.
@@ Methods
To explore the relationship between crime and commuting, we use police-recorded crime data from 2011/2012 to match the census period. Recording practices vary across police forces and between countries in the UK. In England and Wales, crime data is attributed to Community Safety Partnerships (CSPs), local collaborations that aim to reduce crime and improve community safety. For our analysis, we use the “Police Recorded Crime Community Safety Partnership Open Data Table, year ending March 2012” data set, which provides detailed records of all recorded offences per CSP from April 2011 to March 2012  (?). Because this data set is more detailed and consistent than those available for other parts of the UK, we focus exclusively on England and Wales, excluding Scotland and Northern Ireland from our crime analysis. Additionally, since some CSPs encompass multiple local authorities—for example, North Yorkshire CSP includes Craven, Hambleton, Harrogate, Richmondshire, Ryedale, Scarborough, and Selby—we pre-processed the data to aggregate nodes from the local authority commuting data network to their corresponding CSPs used in the 2011 crime data set. To associate each LAD with its respective CSP, we use the “LAD to Community Safety Partnership to PFA (December 2019)” data set (?), which provides the most consistent structural match with the 2011 Census commuting data. The “UK Local Authorities (past and current)” data set (?) was used to obtain LAD codes—essential for linking LADs to CSPs—and to retrieve geographic coordinates for constructing the England and Wales network. Finally, we used the “2011 Census: Usual resident population, local authorities in England and Wales” to obtain the estimated total population for each LAD (?), which we preprocessed by aggregating multiple LADs into their corresponding CSPs and summing their populations where CSPs spanned more than one LAD.
@@
For burglary, Model 2 offers the best fit based on BIC and includes only statistically significant predictors (Table 1). For theft, Model 3 slightly outperforms Model 2 in BIC but only marginally; given the minimal gain in explanatory power and added complexity, Model 2 remains the most robust and interpretable choice (Table 2). Supplementary Note 1 provides the tables for the remaining offenses.
We note that resident population and inbound commuting are correlated across cities, but diagnostics show that this does not pose a problem of multicollinearity in Model 2. The condition number of the design matrix is $158$ and the variance inflation factor (VIF) for the regressors is $1.96$, values that indicate moderate correlation but no regression instability and negligible variance inflation. As an additional robustness check, we re-estimated the model using ridge regression, with the penalty parameter $\lambda$ selected by leave-one-out cross-validation (see Supplementary Note 2), following the procedure of previous studies (?, ?, ?). We find that the ridge regression coefficients are very similar to the OLS estimates. For example, for theft, we find with ridge regression $\gamma=0.91$ (95% CI [0.80, 1.01]) and $\alpha=0.32$ (95% CI [0.22, 0.42]), whereas OLS estimates are $\gamma=0.92$ (95% CI [0.83, 1.01]) and $\alpha=0.32$ (95% CI [0.25, 0.39]), and, for burglary, we find with ridge regression $\gamma=1.05$ (95% CI [0.96, 1.15]) and $\alpha=0.21$ (95% CI [0.14, 0.27]), whereas OLS estimates are $\gamma=1.06$ (95% CI [0.97, 1.15]) and $\alpha=0.20$ (95% CI [0.14, 0.27]). This similarity is also the case for all the other offenses and demonstrates that multicollinearity does not affect coefficient stability in our analysis, with ridge estimates being just slightly smaller than OLS, as expected due to penalization.
@@
BIC 328.3-328.3 357.6-357.6 354.9-354.9
F-statistic 1316.91316.9*** 750.8750.8*** 504.7504.7***
Durbin-Watson $2.147$ $2.168$ $2.159$
Cond. No. $121$ $158$ $6468$